OpenStudy (anonymous):

circle, center (-4,2) radius of 3 what is the equation?

Directrix (directrix):

In the standard form of a circle equation, (h,k) are the coordinates of the center and r is the radius. Post what you get for the answer and someone will check your work, if you would like that. @dancingolivia

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=9?

Directrix (directrix):

Not quite.

Directrix (directrix):

@dancingolivia >>(x-h)^2+(y-k)^2=9? ---> You have to fill in for the center. (x - (-4))^2 + (y - 2)^2 = 3^2 (x + 4)^2 + (y-2)^2 = 3^2